# XpookyNet: Advancement in Quantum System Analysis through Convolutional Neural Networks for Detection of Entanglement

### v3

@@ III.2 Training CNNs
IV Results Evaluation
Figure 6: The proposed XpookyNet performance in the learning process. (a) Accuracy, and (b) Loss function graphs for the methods listed
Table 1: Overview of Methods used in this study for two-qubit entanglement detection. Conv., Brch., Sep., and Plat. stand for convolutional, branched model, separable convolution, and ReduceLROnPlateau, respectively.
Model Structure
Methods
ACC
FNR
MCC
Epoch Time
Number of Conv. Layer
Number of Parameters
NN
0.8325
0.1826
0.6603
19s44ms
0
3,233
Simple Conv.
0.9500
0.0622
0.9189
51s55ms
10
1,334,913
Brch.
0.9624
0.377
0.9276
63s22ms
14
2,015,521
BN. Sep.
0.9632
0.0504
0.9205
57s20ms
10
1,080,263
Plat.
0.9789
0.0244
0.9576
48s33ms
10
1,334,913
@@ Methods
Plat. BN. Sep.
0.9749
0.0440
0.9452
60s47ms
10
1,080,263
Brch. BN. Sep.
0.9664
0.0518
0.9315
75s62ms
14
1,324,615
Brch. Plat. BN. Sep.
0.9761
0.0440
0.9452
75s95ms
14
1,335,430
We assess XpookyNet for detecting entanglement in quantum systems. Our analysis includes loss function and accuracy graphs, overview tables, and classification performance using relevant metrics:
$\small ACC=\frac{TP+TN}{TP+TN+FP+FN}$ (15)
$\small FNR=\frac{FN}{TP+FN}$ (16)
$\small MCC=\frac{TP\times TN-FP\times FN}{\sqrt{(TP+FP)(TP+FN)(TN+FP)(TN+FN)}}$ (17)
IV.1 Two-qubit Entanglement Detection
The NN is a baseline with the lowest ACC, MCC, and a notable FNR. Simple Convolution improves ACC and MCC but retains a discernible FNR. The Branched CNN enhances accuracy but results in a higher FNR than Simple Convolution, albeit with an improved MCC. BN and Separable convolution layers achieve a commendable balance with high ACC, a lower FNR, and an impressive MCC. The Reducing the learning rate on plateaus (ReduceLROnPlateau) outperforms all others with the highest ACC, MCC, and minimal FNR. Combining branched CNN with the ReduceLROnPlateau method is the most optimal choice.
In only 14 epochs, XpookyNet (”Brch.Plat.”) achieves a 98.53% accuracy. The network can even determine the entanglement degree with mean absolute error 0.0034. Graphs 6 (a) depict the performance evolution of methods over epochs. The NN starts with high accuracy but flats around 0.8 ACC. In comparison, convolutional models show progressive improvement, especially the ones that exploit the ReduceLROnPlateau technique. In Graph 6 (a), the combination of the branched model with the ”ReduceLROnPlateau” configuration (”Brch.Plat.”) stands out for its significant improvement over time. Graphs 6 (b) illustrate the loss function of the methods to showcase the effect of ReduceLROnPlateau on epochs $8$ and $12$, emphasizing its action on the loss function’s plateau.
@@ IV.1 Two-qubit Entanglement Detection
Figure 7: Based on purity values of 100%, 83%, 56%, and 37%, confusion matrices illustrate the classification performance of the CNN model in three-qubit quantum systems.
IV.2 Three-qubit Entanglement Detection
XpookyNet utilizes separable convolution layers and BNs in a three-qubit setting for enhanced performance on larger tensors compared to traditional convolutional layers.
@@ IV.2 Three-qubit Entanglement Detection
Improving XpookyNet becomes essential, especially for analyzing quantum states at different purity levels to address noise, errors, and imperfections in the NISQ era.
@@
Classes
Purity
ACC
FNR
MCC
Sep.
1
0.9980
0.0000
0.9938
0.83
0.9951
0.0035
0.9848
0.56
0.9731
0.0275
0.9197
0.37
0.9191
0.0605
0.7616
$AB|C$
1
0.9994
0.0030
0.9981
0.83
0.9986
0.0060
0.9956
0.56
0.9888
0.0405
0.9648
0.37
0.9652
0.1395
0.8889
$A|BC$
1
0.9994
0.0030
0.9981
0.83
0.9987
0.0050
0.9959
0.56
0.9932
0.0244
0.9784
0.37
0.9722
0.1090
0.9117
$AC|B$
1
0.9991
0.0035
0.9972
0.83
0.9983
0.0076
0.9946
0.56
0.9900
0.0400
0.9685
0.37
0.9718
0.1110
0.9104
$ABC$
1
0.9997
0.0015
0.9991
0.83
0.9987
0.0045
0.9959
0.56
0.9938
0.0205
0.9806
0.37
0.9795
0.0605
0.9356
Overall 1 0.9978 0.0022 0.9972
0.83
0.9947
0.0053
0.9934
0.56
0.9694
0.0306
0.9618
0.37
0.9039
0.0961
0.8799
XpookyNet (”Brch. Plat. BN. Sep.”) classifies three-qubit systems with high accuracy on density matrices with purity 1, but its performance declines as purity decreases, depicted in Fig. 7. Owing to imperfections and noise, real-world quantum systems lose their entanglement as purity drops. Thus, the purity plunge primarily affects XpookyNet’s classification of partial entangled states that encompass characteristics of separable and entangled states, with an average 11.7% performance shrinkage. As purity decreases, fully entangled states can appear partial entangled or separable, leading to misclassification in XpookyNet. Mixing density matrices with random ones can also blur class boundaries. However, XpookyNet performs strongly, with only a 5.98% performance decrease on average, despite purity drop effects in these two classes. This behavior of XpookyNet aligns with the physical intuition about quantum entanglement and the effects of system imperfections, demonstrating the network’s ability to capture the complex dynamics of quantum systems despite noise and imperfections [45].
XpoolyNet’s performance declines across all classes as purity decreases from 1 to 0.37. Table 2 expresses ACC and MCC decrease while FNR increases, indicating misclassifications. For separable states, purity decline results in a 7.9% ACC decrease and a 23.4% MCC drop. The FNR increases by 0.0605, a notable rise in errors, especially FP. TN and TP remain relatively accurate, pointing to entanglement features obscuring results. Partial entangled states ($AB|C$, $A|BC$, $AC|B$) show less drastic ACC and MCC decrease than separable states. However, FNR increases by an average of 62.5%, indicating significant errors at decimated purity. Fully Entangled (ABC) states experienced a negligible impact on purity decrease, with only a 1.83% ACC drop and a 6.36% MCC decrease. Oddly, three-qubit full entanglement detection functions better than two-qubit problem due to the limited known tripartite entangled set of GHZ, W, and graph states.
V Conclusion and Discussion
We propose XpookyNet, a novel approach using CNNs to detect entanglement in quantum systems. Our model represents density matrices as 3D tensors and employs parallel layers with expanded kernel sizes for improved efficiency. Throughout our experimentation, we systematically evaluate different model architectures and configurations. The results demonstrate that XpookyNet outperforms its counterparts regarding evaluation metrics, effectively detecting entanglement quantum systems. Our analysis also extends to the impact of purity levels on XpookyNet’s prowess. The model’s mastery subsides as purity drops, although it performs flawlessly on high-purity matrices. Nevertheless, in tandem with classifying the quantum system’s conditions, the growth in classes’ errors shows alignment with physical intuition about quantum entanglement, confirming XpookyNet’s robustness and connoisseur of quantum idiosyncrasies. By entrenching a procedure that generates all Hilbert space’s partial entangled states $AB|C$, $AC|B$, and $BC|A$ and a deft model that captures the quantum system’s traits through its density matrix, the detection of tripartite entanglement ABC becomes a cinch since recognizing that the density matrix of a fully entangled $ABC$ system encapsulates the integrals derived from the partial entangled states $AB|C$, $AC|B$, and $BC|A$. By analogy, more qubit entanglement configurations are paved[46, 47]. Moreover, XpookyNet can be considered a criterion for quantum systems due to its high accuracy and understanding of quantum features.
Acknowledgements.
We extend our heartfelt appreciation to the University of Tehran for fostering a collaborative and nurturing environment that greatly benefited the researchers engaged in this study.
@@ B.2 Partial entangled density matrix
Case 1: Partial entangled pairs A and B are generated by:
@@
Case 2: Partial entangled pairs B and C are generated by:
@@
Case 3: Partial entangled pairs A and C are generated by state vectors of the entangled state $|\psi_{AC}\rangle$ and separated state $|\psi_{B}\rangle$ considered as:
@@
$\begin{split}|\psi_{AC|B}\rangle=[a_{0}b_{0}\ \ a_{1}b_{0}\ \ a_{0}b_{1}\ \ a_{1}b_{1}\\ a_{2}b_{0}\ \ a_{3}b_{0}\ \ a_{2}b_{1}\ \ a_{3}b_{1}]^{T},\end{split}$ (25)
where $a$ and $b$ are corespondent magnitudes of equations (23) and (24). Ultimately, the partial entangled states are generated by:
@@ B.5 Three-qubit Graph state
$\left|\Psi_{\text{Graph}}\right\rangle=$ $\frac{1}{\sqrt{N_{\text{Graph}}}}(\alpha_{0}\ |000\rangle+\alpha_{1}\ |001\rangle$
$\quad+\alpha_{2}\ |010\rangle-\alpha_{3}\ |011\rangle+\alpha_{4}\ |100\rangle$
$\quad+\alpha_{5}\ |101\rangle-\alpha_{6}\ |110\rangle+\alpha_{7}\ |111\rangle),$ (33)

### v4

@@ 3.2 Training CNNs
4 Results Evaluation
We assess XpookyNet with different structures for detecting entanglement in quantum systems. Our analysis includes absolute error and accuracy plots, overview tables, and classification and regression performance using relevant metrics.
$ACC=\frac{TP+TN}{TP+TN+FP+FN}$ (16)
Mean absolute error (MAE) measures the average magnitude of errors between predicted values and actual values to asses the regression performance.
$MAE=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\left|y_{i}-\hat{y}_{i}\right|$ (17)
where $N_{s}$ is the total number of instances. $y_{i}$ represents the actual value of the $i^{th}$ instance. $\hat{y}_{i}$ represents the predicted value for the $i^{th}$ instance, and $y_{i}$ is the actual value to asses the regression performance.
$FNR=\frac{FN}{TP+FN}$ (18)
$MCC=\frac{TP\times TN-FP\times FN}{\sqrt{(TP+FP)(TP+FN)(TN+FP)(TN+FN)}}$ (19)
4.1 Two-qubit Entanglement Detection
Table 1: Overview of XpookyNet’s performance with different structures used in this study for two-qubit entanglement detection.
Model detail
Structure ACC FNR MCC Epoch time Number of Conv. layers Number of parameters
NN 0.8325 0.1826 0.6603 19s44ms 0 3,233
Simple Conv. 0.9500 0.0622 0.9189 51s55ms 10 1,334,913
Brch. 0.9624 0.0377 0.9276 63s22ms 14 2,015,521
B.N. Sep. 0.9632 0.0504 0.9205 57s20ms 10 1,080,263
Plat. 0.9789 0.0244 0.9576 48s33ms 10 1,334,913
@@ 4.1 Two-qubit Entanglement Detection
Plat. B.N. Sep. 0.9749 0.0440 0.9452 60s47ms 10 1,080,263
Brch. B.N. Sep. 0.9664 0.0518 0.9315 75s62ms 14 1,324,615
Brch. Plat. B.N. Sep. 0.9761 0.0440 0.9452 75s95ms 14 1,335,430
@@
The NN is a baseline with the lowest ACC, MCC, and a notable FNR. The structure Simple Conv. improves ACC and MCC but retains a discernible FNR. Three branch layers (Brch.) enhances accuracy but results in a higher FNR than Simple Conv., albeit with an improved MCC. B.N. Sep. achieve a commendable balance with high ACC, a lower FNR, and an impressive MCC. The Reducing the learning rate on plateaus (Plat.) outperforms all others with the highest ACC, MCC, and minimal FNR. Combining branched CNN with the Plat. method (Brch. Plat.) is the most optimal structure.
In only 14 epochs, XpookyNet achieves 98.53% accuracy. The network can even determine the entanglement degree EoF with a MAE of 0.0095. Fig. 6(a) depict the performance evolution of methods over epochs. The NN starts with high accuracy but flats around ACC $\approx 0.8$. In comparison, convolutional models show progressive improvement, especially the ones that exploit Plat. technique. Altogether, the combined structure Brch. Plat. stands out for its significant improvement over time. Besides, the effect of Plat. method on epochs $8$ and $12$ is showcased. Fig. 6(b) also compares the decrease of MAE under each method and their combination, pointing out how combining the methods leads to a lower error in estimating the EoF of the states.
Figure 6: The proposed XpookyNet’s performance in the learning process. The progression of each XpookyNet structure’s performance metrics on (a) the classification, and (b) the regression tasks.
4.2 Incomplete Density Matrix
Quantum tomography is a procedure used to determine the complete quantum state of a system. The primary outcome of tomography is the density matrix, which captures a quantum system’s statistical properties and provides a full description of the quantum state. However, tomography is notoriously costly in terms of time and resources due to the exponential growth in complexity with the size of the system [43].
In this subsection, we explore handling incomplete density matrices to evaluate model performance on fewer measurements. We consider the primary dataset described in subsection 2.3 as a training dataset, generated from complete density matrices with full tomography. By relying only on machine learning capacity, this study investigates the possible elimination of the need for costly full tomography, evaluating the performance of XpookyNet in classifying incomplete data.
In quantum tomography, the density matrix is reconstructed by preparing multiple identical copies of the quantum state, in which measurements are repeated on multiple bases that span the whole state space. The measurement operators, which are the basis for an $N$-qubit system, are constructed as tensor products of Pauli matrices ($\sigma_{1},\sigma_{2},\sigma_{3}$), along with the identity matrix ($\sigma_{0}$), resulting in a set of $2^{N}\times 2^{N}=4^{N}$ different operators. The linear combination of these operators assembles the density matrix as follows:
$\rho=\frac{1}{2^{N}}\sum_{\bar{v}}\mathrm{Tr}(\sigma_{v_{1}}\otimes\sigma_{v_{2}}\otimes\cdots\otimes\sigma_{v_{N}}\rho)\sigma_{v_{1}}\otimes\sigma_{v_{2}}\otimes\cdots\sigma_{v_{N}}$ (20)
The summation encompasses vectors $\bar{v}=(v_{1},\cdots,v_{N})$, where each entry $v_{i}$ is selected from the set $i=0,1,2,3$. By conducting measurements of observables that are products of Pauli matrices, we can estimate each component within this summation, thereby enabling us to approximate $\rho$ [44].
In evaluation process, we generate incomplete data from complete data so that XpookyNet’s performance can be assessed using fewer measurements. While a complete density matrix $\rho$ includes all measurement bases $\bar{v}$, an incomplete density matrix comprises only a subset of the total bases. Assuming $\bar{v^{\prime}}$ are the arbitrary ignored measurement bases in the quantum tomography process, the incomplete density matrix $\rho^{\prime}$ can be generated from the complete density matrix as follows:
$\rho^{{}^{\prime}}=\rho-\frac{1}{\mathcal{N}}\sum_{\bar{v^{\prime}}}\mathrm{Tr}(\sigma_{v^{\prime}_{1}}\otimes\sigma_{v^{\prime}_{2}}\otimes\cdots\otimes\sigma_{v^{\prime}_{N}}\rho)\sigma_{v^{\prime}_{1}}\otimes\sigma_{v^{\prime}_{2}}\otimes\cdots\sigma_{v^{\prime}_{N}}$ (21)
In a 2-qubit system, we evaluate 10,000 instances of incomplete test data generated from complete test data according to equation (21). Incomplete data are characterized by a random subset of measurement bases, with varying numbers of ignored measurement bases.
As depicted in Fig. 7, XpookyNet’s performance is a trade-off for the number of measurements with accuracy, differing between entire and Bell entangled spaces. In the entire entangled space, the model maintains around 80% accuracy with half the density matrix elements. Detecting Bell entangled states with fewer measurements is crucial for efficient verification, enhancing the feasibility and scalability of quantum information processing. In Bell entangled space, accuracy remains above 90% with seven measurements and above 80% with only three measurements.
Figure 7: Accuracy plot of XpookyNet’s performance for the classification task on incomplete data in two different data spaces: Entire entangled space and Bell entangled space (The dashed line represents the best model accuracy in Table 1 on complete data: ACC = 0.9852).
4.3 Three-qubit Entanglement Detection
@@
Improving XpookyNet becomes essential, especially for analyzing quantum states at different purity levels to address noise, errors, and imperfections in the NISQ era. Hence, XpookyNet utilizes B.N. Sep. method in a three-qubit setting for enhanced performance on larger tensors compared to traditional convolutional layers.
XpookyNet (”Brch. Plat. B.N. Sep. structure”) classifies three-qubit systems with high accuracy on density matrices with pure state $p=1$, but its performance declines as purity decreases, depicted in Fig. 8. Real-world quantum systems yield less coherence due to imperfections and noise; decoherent systems induce decreased density matrix purity. Since decoherence disrupts quantum correlations within the quantum system, quantum systems lose entanglement as purity reduces [45]. Hence, the purity plunge primarily turns down XpookyNet’s functionality in partial entangled states encompassing the characteristics of both separable and entangled states, with an average of 11.7%. As purity decreases, fully entangled states can appear partial entangled or separable, leading to misclassification in XpookyNet. Mixing density matrices can also blur class boundaries. However, XpookyNet performs strongly, with only a 5.98% performance decrease on average, despite purity drop effects in these two classes. This behavior of XpookyNet aligns with the physical intuition about quantum entanglement and the effects of system imperfections, demonstrating the network’s ability to capture the complex dynamics of quantum systems despite noise and imperfections [46].
Figure 8: Based on purity values of $p=1,0.83,0.56,0.37$, (a) confusion matrices, (b) accuracy plot, and (c) matthews correlation coefficient plot illustrate the classification performance of XpookyNet in three-qubit quantum systems.
@@
ACC FNR MCC
Class/p 1 0.83 0.56 0.37 1 0.83 0.56 0.37 1 0.83 0.56 0.37
Sep. 0.99 0.99 0.97 0.91 0.00 0.00 0.02 0.06 0.99 0.98 0.91 0.76
$AB|C$ 0.99 0.99 0.98 0.96 0.00 0.00 0.04 0.13 0.99 0.99 0.96 0.88
$A|BC$ 0.99 0.99 0.99 0.97 0.00 0.00 0.02 0.10 0.99 0.99 0.97 0.91
$AC|B$ 0.99 0.99 0.99 0.97 0.00 0.00 0.04 0.11 0.99 0.99 0.96 0.91
$ABC$ 0.99 0.99 0.99 0.97 0.00 0.00 0.02 0.06 0.99 0.99 0.98 0.93
overall 0.99 0.99 0.96 0.90 0.00 0.00 0.03 0.09 0.99 0.99 0.96 0.87
XpoolyNet’s performance declines across all classes as purity decreases from 1 to 0.37. Table 2 expresses ACC and MCC decrease while FNR increases, indicating misclassifications. For separable states, purity decline results in a 7.9% ACC decrease and a 23.4% MCC drop. The FNR increases by 0.0605, a notable rise in errors, especially FP. TN and TP remain relatively accurate, pointing to entanglement features obscuring results. As illustrated in Fig. 8, partial entangled states ($|\psi_{AB|C}\rangle$, $|\psi_{A|BC}\rangle$, $|\psi_{AC|B}\rangle$) show less drastic ACC and MCC decrease than separable states. However, the FNR increases by an average of 62.5%, indicating significant errors at decimated purity. Fully Entangled ($|\psi_{ABC}\rangle$) states experienced a negligible impact on purity decrease, with only a 1.83% ACC drop and a 6.36% MCC decrease. Oddly, three-qubit full entanglement detection functions better than two-qubit problem due to the limited known tripartite entangled set of GHZ, W, and graph states.
5 Discussion and Conclusion
In the quest to advance quantum computing, entanglement detection stands as a challenge. Numerous researchers have embarked on this endeavour, employing network-based ML techniques to achieve higher accuracy and productivity. Several contributions are compared here, highlighting their unique methodologies and findings and showing how ours advances them.
M. Yosefpor et al. leverage autoencoder NNs to optimize entanglement detection by identifying a minimal set of measurements for detecting entangled states, reducing experimental costs while achieving the high accuracy of 98% with full tomography [47]. The research by Y. Chen et al.’s method for entanglement detection uses unsupervised learning [48]. Their approach leverages the convexity of non-entangled samples to design a complex-valued network, combining a pseudo-siamese network with a generative adversarial network. This methodology, trained with separable states, achieves a detection accuracy of over 97.5%. N. Asif et al. [49] developed an NN model aimed at classifying quantum states by optimizing a Bell-type inequality for the relative entropy of coherence. Their method, which labels quantum states using the PPT criterion, achieves an accuracy of up to 94.62% and demonstrates scalability to more complex systems. YD. Qu et al. [50] focus on using complex-valued CNN to classify quantum states, achieving an average accuracy of 98.7%, with peak performance at 99.99% for only four-qubit pure states. The researchers emphasize the importance of incorporating phase and amplitude information in datasets to classify accurately. J. Pawlowski et al. [51] uniquely address the identification of positive-under-partial-transposition entangled states (PPTES) through entanglement-preserving symmetry operations within a triple Siamese network. This semi-supervised approach excels in analyzing specific state classes, including Acin, UPB, Horodecki, mixed, and pure states, demonstrating a high accuracy, particularly in detecting pure states.
There are, however, difficulties with generalizability due to their reliance on limited states for training, as well as the computational complexity of their architectures, particularly when dealing with real-world quantum systems. Due to the complexity of their models or their sensitivity to boundary data, scaling up to larger quantum systems requires substantial resources and expertise.
The proposed XpookyNet framework offers a streamlined and efficient approach by representing density matrices as 3D tensors for entanglement detection using CNNs. Compared with its counterparts, XpookyNet has a much higher level of accuracy in terms of pervasive evaluation metrics, as it inherits parallel layers and utilizes tailored kernel sizes; it learns data quickly with fewer training epochs due to its devised confrontation with plateauing.
The versatility of XpookyNet allows it to be used for regression tasks to estimate the EoF of a density matrix. Additionally, our analysis also extends to the impact of decoherence on fully and partially entangled states, considering the purity of density matrices. The model’s mastery subsidence with purity drop, in tandem with the growth in classes’ errors, shows alignment with physical intuition about quantum entanglement, confirming XpookyNet’s connoisseur of quantum idiosyncrasies. By establishing a procedure that generates all of Hilbert space’s partial entangled states $|\psi_{AB|C}\rangle$, $|\psi_{AC|B}\rangle$, and $|\psi_{BC|A}\rangle$ and employing a model that accurately captures the quantum system’s characteristics through its density matrix, detecting tripartite entanglement $|\psi_{ABC}\rangle$ becomes straightforward as a multi-classification task. This is because the density matrix of a fully entangled $|\psi_{ABC}\rangle$ system encapsulates the integrals derived from the partial entangled states $|\psi_{AB|C}\rangle$, $|\psi_{AC|B}\rangle$, and $|\psi_{BC|A}\rangle$, paving the way for more qubit entanglement configurations by analogy [52, 53]. Importantly, our research devises a method to assess the model’s performance with fewer measurements, striking a balance between accuracy and the cost of tomography.
[…]
Acknowledgements No acknowledgments to declare.
[…]
Contribution Statement AK: Conceptualization, Methodology, Investigation, Programming, Literature survey, Writing. YM: Data Adjustment, Visualization, Literature Survey, Writing. PK: Literature survey, Revising Manuscript. HA: Revising Manuscript, Project administration, supervision. KF: Revising Manuscript, Supervision. MB: Revising Manuscript, Supervision.
[…]
Funding
Not applicable.
[…]
Conflict of interest
Not applicable.
[…]
Codes availability
The codes that support the plots within the paper are publicly available on: https://github.com/AKookani/XpookyNet
@@ B.2 Partial entangled density matrix
Case 1: Partial entangled pairs $|\psi_{A}\rangle$ and $|\psi_{B}\rangle$ are generated by:
@@
Case 2: Partial entangled pairs $|\psi_{B}\rangle$ and $|\psi_{C}\rangle$ are generated by:
@@
Case 3: Partial entangled pairs $|\psi_{A}\rangle$ and $|\psi_{C}\rangle$ are generated by state vectors of the entangled state $|\psi_{AC}\rangle$ and separated state $|\psi_{B}\rangle$ considered as:
@@
$|\psi_{AC|B}\rangle=[a_{0}b_{0}\ \ a_{1}b_{0}\ \ a_{0}b_{1}\ \ a_{1}b_{1}\ \ a_{2}b_{0}\ \ a_{3}b_{0}\ \ a_{2}b_{1}\ \ a_{3}b_{1}]^{T},$ (29)
where $a$ and $b$ are corespondent amplitudes of states (27) and (28). Ultimately, the partial entangled states are generated by:
@@ B.5 Three-qubit Graph state
$\left|\Psi_{\text{Graph}}\right\rangle=$ $\frac{1}{\sqrt{N_{\text{Graph}}}}(\alpha_{0}\ |000\rangle+\alpha_{1}\ |001\rangle+\alpha_{2}\ |010\rangle-\alpha_{3}\ |011\rangle$
$\quad+\alpha_{4}\ |100\rangle+\alpha_{5}\ |101\rangle-\alpha_{6}\ |110\rangle+\alpha_{7}\ |111\rangle)$